Tr3:

Michael bikes 36 miles last week and 38.4 miles this week. What is the average distance he bikes? Give your answer with the appropriate precision. A. 36 mi B. 36.2 mi C. 37 mi D. 37.2 mi

Ultrilliam:

@Vocaloid

Ultrilliam:

add the 2 values together, 36+38.4 then divide them by how many there are, in this instance 2. so (36+38.4)/2=?

Ultrilliam:

@Tr3

Tr3:

thx

Vocaloid:

Make sure to consider significant figures (since the question asks)

Vocaloid:

Since your smallest number of sig figs is 2 then your answer should Also have two dog figs

Vocaloid:

*sig figs

Vocaloid:

In this particular situation your final answer would be a two digit whole number (no decimal)

Allison:

36+38.4 = 74.4/2 = D. 37.2 ._.

Vocaloid:

Give your answer with the appropriate precision." This implies that significant figures are taken into account, so the final answer would have 2 significant figures. 37.2 has 3 significant figures whereas the answer must have exactly 2 significant figures based on the number 36 from the original problem

Allison:

So it's C?

Tr3:

idk